# Turn Every Application into an Agent: Towards Efficient Human-Agent-Computer Interaction with API-First LLM-Based Agents

### v1

@@ 5.2. Participants
5.3. Subjective Metric Collection
As mentioned in 5.1, we used four questionnaires to collect users’ subjective experence for completing the tasks using different methods. Questionnaires 1 to 4 were administered separately after completing L1 tasks manually, completing L2 tasks manually, completing L1 tasks with the assistance of an Agent, and completing L2 tasks with the assistance of an Agent, respectively.
To determine whether LLM-based agents can indeed reduce users’ cognitive load for completing tasks compared to manual work, in all the four questionnaires, we include questions based on the NASA Task Load Index (NASA-TLX) (Hart, 1988) and an additional question on the required learning efforts. For NASA-TLX, our questions cover all the six metrics, including Mental Demand, Physical Demand, Temporal Demand (how hurried or rushed of the tasks), performance (feeling of success in completing the task), frustration level, and completion effort (how hard the users need to work on the tasks). Lower values in those six metrics indicate higher cognitive loads, better feeling of success, lower frustration level, and less efforts during task completion. For the learning efforts, lower score also indicates less learning efforts.
To further compare users’ perceptions on the UI Agent and AXIS, Questionnaires 3 and 4 contained questions on the ratings on fluency, reliability, UI dependency, decision consistency, and perceived speed for both Agents. Specifically, UI dependency measures the degree of users’ reliance on the UI while observing the Agent complete tasks. Decision consistency assesses how closely the decisions made by the Agent align with the decisions users might make to complete the same tasks manually (the experimental web page displayed all the decisions made by the Agents in a step-by-step manner). Perceived speed refers to users’ subjective perception on how fast the Agents completed the tasks.
5.4. Objective Metric Collection
We recorded experimental logs throughout the experiment, including screen recordings of the manual completion of tasks by the users, screen recordings of the two Agents (UI agent and AIXS) performing tasks, decision-making processes, UI interaction paths, time taken, and the cost of the LLM backend (GPT-4o, version 20240513). From on the logs, we extracted objective metrics including the time and success rate of task completion across the three methods, the degree of UI dependency, and cost for the two Agents.
6. Results
In line with our research questions, we divided the experimental results into three parts for analysis. Firstly, We investigate how the adoption of agent reduces cognitive load for users. Secondly, We analyze and compare different agents’ abilities of enhancing task efficiency. Finally, we discuss users’ preferences between the traditional UI agent and our AXIS agent.
6.1. Cognitive load
To investigate the reduction of cognitive load by LLM-based Agents, we analyzed the NASA-TLX and learning effort scoring collected from users during the task execution process and summarized results in Table 6 and Figure 3. It is worthy noting that, in our experiment, L2 tasks generally scores higher than L1 tasks across multiple dimensions of the NASA-TLX scale and learning efforts, which indicates that our task difficulty classification is consistent with the users’ experience.
As shown in Table 6, at both L1 and L2 difficulty levels, the agent based method shows significant improvements over the manual based method in most of the NASA-TLX scales. In particular, the agent based method is much less mentally and physically demanding, generate less frustration for users, and requires less effort in completing tasks than the manual based method. The reduction in the cognitive effort by the agent based method is also generally more pronounced for the more difficulty L2 tasks. For the performance metric, while the difference between agent based method and manual method is insignficant for the easy L1 tasks, the agent based method does boost the users’ feeling of success significantly (p¡0.05) for the difficult L2 tasks. In Figure 3 (b), we also summarized the average scores across all the six NASA-TLX scales. This figure shows that, when using the Agents, the users’ experiences in completing L1 and L2 tasks are similar, which demonstrates the stability of LLM-based agents in addressing tasks with different complexities. Finally, Figure 3 (c) shows that the LLM-based agents can also significantly reduced users’ learning efforts in completing the task compared to the manual approach. Similar to the NASA-TLX scales, the reduction is also bigger at the higher task difficulty level.
All those results clearly demonstrate the value of agent based method in helping users in completing various tasks and answer the first research question (RQ1): the LLM-based agent does indeed lower the cognitive load of users and reduces their effort to learn, especially for more difficult tasks.
| Metric | Task Level | Manual | Agents | Pairwise Significance |
| Mental Demand (0-100) | L1 L2 | 21.3 70.0 | 2.5 7.5 | L1: m¿a (p ¡ 0.001) L2: m¿a (p ¡ 0.001) |
| Physical Demand (0-100) | L1 L2 | 31.3 57.5 | 5.0 6.3 | L1: m¿a (p ¡ 0.001) L2: m¿a (p ¡ 0.001) |
| Temporal Demand (0-100) | L1 L2 | 52.5 37.5 | 28.8 35.0 | L1: m¿a (p ¡ 0.05) L2: - |
| Performance (0-100) | L1 L2 | 21.2 47.5 | 21.2 26.2 | L1: - L2: m¿a (p ¡ 0.05) |
| Frustration Level (0-100) | L1 L2 | 31.3 62.5 | 7.5 10.0 | L1: m¿a (p ¡ 0.001) L2: m¿a (p ¡ 0.001) |
| Completion Effort (0-100) | L1 L2 | 12.5 35.0 | 17.5 13.8 | L1: - L2: m¿a (p ¡ 0.01) |
6.2. Efficiency and reliability
To compare the efficiency and reliability between the manual method, UI agents, and AXIS, we also collected metrics on the completion time, success rate, as well as the number of steps and costs for completing tasks in our experiment. Those information are summarized in Table 7 and Table 8.
In terms of time efficiency, AXIS consistently took significantly less time than both the manual and UI Agent methods for both L1 and L2 tasks (p ¡ 0.001), with a larger advantage for the more difficult L2 tasks. For L1 tasks, the manual method was actually faster than the UI Agent as the UI Agent generally took many steps to complete a task.
For the accuracy, unsurprisingly the manual method is the best among all the three methods. Still, AXIS can achieve a high level accuracy that is only slightly worse than human performance. In contrast, the accuracy of the UI agent is considerably lower, especially at the L2 tasks. Upon reviewing the video logs, we found that incorrect positioning of UI elements and the invisibility of certain UI components were the major causes of the errors made by the the UI agent.
Finally, as shown in Table 8), UI agents often need to take more steps to finish the task, especially the L2 tasks where where the target UI elements were buried deeper in the interface. In contrast, empowered by the streamlined task execution enabled by API calls, AXIS took significantly fewer steps, and thus lower overall costs for completing tasks at both difficulty levels.
Based on the above analysis, we can address our second research question (RQ2). The use of a UI Agent provides a slight improvement in human efficiency for specific complex tasks, but it suffers from reliability issues. In contrast, AXIS consistently enhances human efficiency and demonstrates greater reliability.
| Metric | Task Level | Manual | UI Agent | AXIS | Pairwise Significance |
| Time(s) | L1 L2 | 61.8 167.6 | 104.6 155.5 | 18.2 57.1 | L1: m¡u (p ¡ 0.001) L1, L2: a¡m (p ¡ 0.001) L1, L2: a¡u (p ¡ 0.001) |
| Success Rate(%) | L1 L2 | 100.0 97.5 | 75.0 45.0 | 98.3 95.0 | L1, L2: m¿u (p ¡ 0.001) L1, L2: a¿u (p ¡ 0.001) |
| Metric | Task Level | UI Agent | AXIS | Pairwise Significance |
| steps | L1 L2 | 6.4 11.1 | 1.0 4.2 | L1: a¡u (p ¡ 0.001) L2: a¡u (p ¡ 0.001) |
| cost($) | L1 L2 | 0.6 0.9 | 0.07 0.3 | L1: a¡u (p ¡ 0.001) L2: a¡u (p ¡ 0.001) |
6.3. Affinity preference
To explore the differences in user experience when executing tasks with UI agents versus AXIS, We also conducted a subjective preference evaluation on five aspects for both L1 and L2 tasks and summarized the results in Figure 4. At both difficulty levels, participants showed a general preference for AXIS over UI agents in term of the perceived speed, fluency, and reliability.
For the perceived consistency of decision, the results varied between the L1 and L2 tasks. In the L1 tasks, AXIS usually can complete tasks with one or a few steps due to the high encapsulation of its API, which resulted in a decision-making style that is quite distant from the thinking pattern of human. However, in complex tasks, AXIS’s decisions become more aligned with human’s thought processes and were thus perceived better by humans in this aspect. For the UI dependency, AXIS is much less reliant to the UI compared to the UI Agent, which needs to frequently interact with the UI interface. This reduced dependency on the UI was clearly perceived by users during their experience.
In summary, for the third research question (RQ3): AXIS, compared to UI agents, tends to leave users with a perception of better efficiency, smoothness, and reliability. As task complexity increases, users are also more inclined to favor AXIS for task resolution. Additionally, feedback from user surveys indicated a desire for greater control when using LLM-based agents. Unlike UI agents, which often interrupt users’ mouse actions and occupy screen space, AXIS’s API-first approach addresses this issue and improves users’ experience.
7. Discussion
7.1. AXIS help to digest unnecessary Application UIs
To illustrate this process, in Figure 5, the UI hierarchical relationships between UIs are represented as a tree, in which each node represents a UI element with higher-level UI elements as parent nodes and lower-level ones as child nodes. We further use red nodes to represent UI locations that can be API-ified after explored by AXIS, and use blue nodes represent general UI elements. In this example, the root node that represents the ”Home” tab is a blue node as not all its sub-UI nodes are red (API-ified). However, the second-level node ”Highlight Color” (node 2-2) and all its third-level child nodes can all be API-ified and are colored in red. Generally, we define a node N as non-essential if this node along with all their child nodes can all be API-ified:
7.2. Turn An Application into an Agent
8. Conclusion
In this paper, we introduce AXIS, a pioneering framework designed to enhance human-agent-computer interaction (HACI) and address the inefficiencies and cognitive burdens associated with multimodal large language models (MLLMs) in complex task execution by prioritizing API calls over traditional UI interactions. Through user experiments with tasks from Office Word , AXIS has proven to be highly effective, reducing task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining a high level of accuracy comparable to human performance. These results underscore the potential of API-first LLM-based agents to streamline interactions, minimize latency, and enhance reliability in task execution.
Our research contributes to the broader field of human-agent interaction by highlighting the limitations of traditional UI-based approaches and proposing a novel solution that leverages API calls to simplify and accelerate task completion. By enabling applications to act as agents through a reduced set of essential UIs and enhanced API sets, AXIS not only improves efficiency but also paves the way towards the development of a comprehensive Agent OS. This paradigm shift suggests that every application has the potential to transform into an intelligent agent capable of executing tasks with minimal user intervention.
In conclusion, AXIS represents a significant step forward in reducing cognitive load and enhancing the efficiency of task completion with LLM-based agents. Our findings provide valuable insights for application developers and researchers, encouraging them to rethink UI designs and explore new ways to integrate API-driven interactions. In our future work, we will focus on extending this framework to a broader range of applications, exploring its impact on various user groups, and advancing the potential of LLMs in creating a more intuitive and efficient human-computer interface.
Appendix A User Study Web Interface
@@
Appendix B User Study Tasks
We sampled five tasks about Microsoft Word in user study which were categorized into low difficulty (L1) and high difficulty (L2). Here are the detailed tasks in table 9.
@@ Appendix B User Study Tasks
| 1 | Here is an article, type in a title ”Impossible Friendship between mouse and cats” and set the title in the center with ”Arial” type of 20 font size. | L1 |
| 2 | Insert a header named ”header” and a footer named ”footer”. | L1 |
@@
Appendix C User Study Survey Form
@@
C.1. Cognitive load related forms
@@
C.2. Human Preference related forms
@@
Appendix D Feasibility Study Tasks
In the feasibility study, we randomly sampled 50 tasks from the WikiHow page ’Use Microsoft Word’ and the official Microsoft Word website. To increase the task difficulty, some of the 50 tasks were composed of smaller sub-tasks, thus increasing the number of steps required for completion. As the tasks sampled in user study, the 50 tasks were also divided into 2 levels of difficulty: low difficulty (L1) and high difficulty (L2), based on factors such as the number of UI interactions required, the depth of the UI functions, and the number of ribbon switches. Table 10 has shown the distribution of the required execution steps (i.e., the number of steps a human would typically need to perform) of the 50 tasks, along with the number of tasks in different difficulty levels.

### v2

@@ C.2 Participants Recruitment
C.3 User Study Web Interface
@@
C.4 User Study Tasks
We sampled five tasks about Microsoft Word in user study which were classified into low difficulty (L1) and high difficulty (L2) following the same criteria as tasks in feasibility study. Here are the detailed tasks in Table LABEL:tab:exp-tasks.
@@ C.4 User Study Tasks
| 1 | Here is an article, type in a title "Impossible Friendship between mouse and cats" and set the title in the center with "Arial" type of 20 font size. | L1 |
| 2 | Insert a header named "header" and a footer named "footer". | L1 |
@@
C.5 User Study Survey Form
@@
C.5.1 Cognitive load related forms
@@
C.5.2 Human Preference related forms
@@
Appendix D Feasibility Study
In the feasibility study, we randomly sampled 50 tasks from the WikiHow page ’Use Microsoft Word’ and the official Microsoft Word website. To increase the task difficulty, some of the 50 tasks were composed of smaller sub-tasks, thus increasing the number of steps required for completion. As the tasks sampled in user study, the 50 tasks were also divided into 2 levels of difficulty: low difficulty (L1) and high difficulty (L2), based on three key factors for task difficulty classification:
Number of UI Interactions Required: Measured by the procedural steps (validated against manuals or user testing), shown quantitatively in Table LABEL:tab:50tasks-steps.
Depth of UI Functions: Evaluated via the Control Hierarchy Tree (rooted at the Ribbon level, depth = 0). For example, the task Apply text highlighting (Home → Text Highlight Color) has a depth = 1.
Ribbon Switches Counts: Refers to cross-tab interactions.
A task is classified as L2 difficulty only if it satisfies all three criteria simultaneously: UI interactions, Maximum control depth and Ribbon switch.
Table LABEL:tab:50tasks-steps has shown the distribution of the required execution steps (i.e., the number of steps a human would typically need to perform through UI) of the 50 tasks, along with the number of tasks in different difficulty levels.
@@ Appendix D Feasibility Study
Appendix E Extensive Applications of AXIS
E.1 AXIS help to digest unnecessary Application UIs
To illustrate this process, in Figure LABEL:fig:lessui, the UI hierarchical relationships between UIs are represented as a tree, in which each node represents a UI element with higher-level UI elements as parent nodes and lower-level ones as child nodes. We further use red nodes to represent UI locations that can be API - ified after explored by AXIS, and use blue nodes to represent general UI elements. Red nodes imply that the UI controls at these locations can be replaced by APIs. These red - marked UI elements can be described using natural language. In contrast, blue nodes indicate that the corresponding controls are either difficult to describe in language or lack corresponding triggering APIs. In this example, the root node that represents the "Home" tab is a blue node as not all its sub-UI nodes are red (API-ified). However, the second-level node "Highlight Color" (node 2-2) and all its third-level child nodes can all be API-ified and are colored in red. Generally, we define a node N as non-essential if this node along with all their child nodes can all be API-ified:
E.2 Turn An Application into an Agent